# Understanding In-context Learning of Addition via Activation Subspaces

### v2

@@ 3.2 Further Refinement via Ablations
We suspect many heads are primarily responsible for storing formatting information (such as ensuring the output appears as a number) rather than encoding information about $k$ itself. Intuitively, while we require the overall signal transmitted by these heads, we do not need any information about the specific value of $k$. To test this hypothesis, we perform mean-ablations: replacing each task-specific signal $h_{k}$ with the overall mean $\overline{h}$ across all values of $k$2.2).
@@
$v=\sum_{h\in\mathcal{H}_{0}}\overline{h}+\sum_{h\in\mathcal{H}_{\text{sig}}\setminus\mathcal{H}_{0}}h_{k}.$ (2)
@@
Focusing on two layers via layer-wise ablation. To efficiently narrow down the important heads, we first perform mean-ablations at the level of layers. From Figure 3(a), we observe that the significant heads $\mathcal{H}_{\text{sig}}$ are concentrated primarily in the middle and late layers. We speculate that heads in the late layers mainly contribute to formatting the output, as they appear too late in the computation to meaningfully interact with the query. After trying different sets of layers, we found that mean-ablating all significant heads outside layers $13$ and $15$ still achieves an intervention accuracy of $0.83$, while mean-ablating any other combination of layers causes negligible drops in accuracy (Appendix A, Table 3). After these ablations, only $11$ heads located in layers $13$ and $15$ remain.
Identifying three final heads via head-ablation. To understand the individual contributions of each head within layers 13 and 15, we perform mean-ablations at the level of individual heads. We first assess the intervention accuracy when retaining only the output of a single head while mean-ablating all other significant heads; however, this generally results in low accuracy. We hypothesize that the output magnitude of a single head is too small to significantly influence the model output, even if it encodes task-relevant information. To amplify each head’s effect, we scale its output by a coefficient (e.g., 5). We find that three heads—head 1 = $(15,2)$, head 2 = $(15,1)$, and head 3 = $(13,6)$—achieve intervention accuracies close to the clean accuracy when appropriately scaled, while all other heads show much lower accuracies regardless of scaling (Table 1). This suggests that these three heads individually encode the task information much better than any others.
@@
[…]
Any other head  Optimal   0.19
@@ 4 Characterizing the Aggregator Subspace
For the model to perform add-k, it has to infer the task information (the number $k$) from the ICL examples. Our next goal is thus to understand how task information is represented in the activation space. Since we have identified three aggregator heads that carry almost all of this information, we can now focus on analyzing the representation space of these three heads.
In this section, we dissect their activation spaces in three stages: (1) Localize a six–dimensional task subspace in each head via principal component analysis (PCA) (§4.1); (2) Rotate this subspace into orthogonal feature directions aligned with sinusoidal patterns across $k$4.2); (3) Decompose the six-dimensional space into a four-dimensional unit-digit subspace and a two-dimensional magnitude subspace that separately encode the units and tens of the answer (§4.3).
@@ 4.1 Localizing to Six-dimensional Subspace
To reduce the $128$-dimensional head activation to a more tractable space to study, we first perform PCA on the $30$ task vectors and find that just six directions can explain $97\%$ of the task variance (Figure 4). We then check that the function vectors found earlier remain effective after projecting onto the subspace. Specifically, we replace each head vector $h_{k}$ with its projection onto the subspace $\tilde{h}_{k}$ to obtain a new function vector $\tilde{v}_{k}=\sum_{h\in\mathcal{H}}\tilde{h}_{k}$ (a variant of $v_{k}=\sum_{h\in\mathcal{H}}h_{k}$ in §2.2). We find that $\tilde{v}_{k}$ has intervention accuracy $0.76$, which is close to the intervention accuracy of $0.79$ before projection (§3.2). Thus, we confine our study to this concise six-dimensional subspace for each head. The rest of the section focuses on head 1 and we defer analogous results for head 2 and 3 to appendix B.
Figure 4: The explained variance ratio vs. number of PCs for each head. The first six PCs make up most of the explained variance ($97\%$) for each head.
Figure 5: Coordinates of head 1’s vectors (their inner products with a PC) for the first six PCs across different add-$k$ tasks. The coordinates in the first five PCs reveal periodic patterns.
@@ 4.2 Identifying Feature Directions Encoding Periodic Patterns
To understand how the six-dimensional subspace of a head represents the task information, we first examine the coordinates of the head vectors (their inner products with the PCs) as a function of $k$. This reveals partially periodic patterns in the first five components (Figure 5).
@@
To find trigonometric functions to fit, we searched over different periods and phases and performed least squares regression. We found six trigonometric functions at periods $2$, $5$, $10$, $10$, $25$, and $50$ that could be expressed as functions of the top 6 PCs with low regression error (Figure 6). We apply the resulting linear transformation to the six PCs to obtain a new set of directions that encode these six pure periodic patterns, which we call feature directions.
Figure 6: The coordinate functions (inner product with feature directions) of head 1’s vectors can fit six trigonometric functions with periods $2,5,10,10,25$, and $50$ well by linear transformation.
@@ 4.3 Decomposing to Subspaces Encoding Subsignals
By construction, the coordinate function of each feature directions (viewed as linear projections of $h_{k}$) is a periodic function of $k$. Mathematically, a feature direction with period $T$ carries task information from the head vectors with “modulo $T$”. Based on this, we hypothesize: (i) the feature direction corresponding to period two, which we call the “parity direction”, encodes the parity of $k$ in the add-k task; (ii) the subspace spanned by the feature directions with periods $2,5,10$, which we call the “unit subspace”, encodes the unit digit of $k$; (iii) the subspace spanned by the directions with periods $25,50$, which we call the “magnitude subspace”, encodes the coarse magnitude (i.e., the tens digit) of $k$.
We verify these hypotheses through causal intervention. In Figure 7, we validate hypothesis (ii): we establish (1) sufficiency by showing that projecting a head vector onto the subspace preserves the relevant task signal; and (2) necessity by showing that projecting a head vector out of the subspace (i.e., onto its orthogonal complement) destroys the relevant task signal. We include analogous experimental results for the two other hypotheses of head 1 in Appendix B.2.
Figure 7: The error rates for the unit digit and the final answer across tasks when projecting head 1’s vectors onto the “unit subspace” (left) and out of the “unit subspace” (right). Projecting onto the unit subspace results in a low unit-digit error rate even when final-answer error remains high, while projecting out leads to high unit-digit error rates that almost fully account for the final-answer errors. This confirms that the unit subspace specifically encodes the unit-digit signal.
5 Signal Extractors of ICL Examples
Previously, we localized the model’s behavior to three heads and their corresponding six-dimensional subspaces, then examined how the model represents the task information ($k$ for add-k) inferred from the ICL examples in one subspace. Now, we analyze how the model extracts the task information from the ICL examples.
In this section, we find that: (1) the signal is primarily gathered from the $y_{i}$ tokens in each example $x_{i}\rightarrow y_{i}$; (2) each example $x_{i}\rightarrow y_{i}$ individually contributes a signal $y_{i}-x_{i}$ in the subspace even on “mixed” in-context examples with conflicting task information; and (3) when all examples $x_{i}\rightarrow y_{i}$ share the same value for $y_{i}-x_{i}$, the extracted signals exhibit a self-correction behavior. For simplicity, we focus on head 1 in this section.
@@ 5.1 Mathematical Observation: Tracing Subspace back to Previous Tokens
Formally, a head $h$’s output at the last token of a prompt $p$ can be written as
$h(p)=\sum_{t\in p}\alpha_{t}\cdot O_{h}V_{h}\cdot z_{t},$ (3)
@@
Let $W_{h}$ denote the projection matrix onto the six-dimensional subspace for head $h$. Then the projected signal at the final token, $W_{h}\cdot h(p)$, can be decomposed into contributions from previous tokens as $W_{h}\cdot\alpha_{t}O_{h}V_{h}z_{t}$, each lying in the transformed head subspace defined by the matrix $W_{h}\cdot O_{h}V_{h}$. In the following subsection, we analyze the magnitudes and directions of these signals, and study how signals from different examples interact.
5.2 Signal Extractor for each Example
Figure 8: For a mixed-$k$ ICL prompt, the strength of the task-signal contribution of each previous token $t$ to the final token, $\|\alpha_{t}W_{h}O_{h}V_{h}z_{t}\|$. Decomposing it into two parts: (1) the norm of the extracted information $\|W_{h}O_{h}V_{h}z_{t}\|$ (top), and (2) the attention score from the final token $\alpha_{t}$ (bottom), both parts consistently peak at the tokens $t=y_{i}$ (in bold).
@@ 5.2 Signal Extractor for each Example
To understand how examples contribute to model generation at the last token, we identify which tokens contribute the most, then examine what information they provide. By the analysis above, the task-signal contribution of each previous token to the final token through the head $h$ is $\alpha_{t}W_{h}O_{h}V_{h}z_{t}$. This can be decomposed into two parts: (1) extracted information: $W_{h}O_{h}V_{h}z_{t}$, the residual stream input projected into the relevant subspace; and (2) aggregation weight: $\alpha_{t}$, the attention score of the final token to the previous token. We plot the norms of the extracted information and the aggregation weights for a random mixed-$k$ ICL prompt in Figure 8. Both the strength of the extracted information and the aggregation weights peak at $y_{i}$ tokens.
@@
We next examine what specific information is extracted from each of these tokens. To do so, we measure the inner product $\langle W_{h}O_{h}V_{h}z_{t},\tilde{h}_{k}\rangle$ between the extracted information and the head vector (projected onto the subspace and normalized to have unit norm) for each task $k$. In Figure 9, we plot this quantity for a random mixed-$k$ ICL prompt for each token $y_{i}$ ($i\in\{1,\ldots,5\}$) and each task $k\in\{1,\ldots,30\}$. We find that the inner product consistently peaks at $k=y_{i}-x_{i}$, indicating that the model extracts the information of $y_{i}-x_{i}$ from the corresponding example $x_{i}\rightarrow y_{i}$.
Figure 9: For a random mixed-$k$ ICL prompt, the inner product between the projected signal from each $y_{i}$ and each head vector (projected onto the subspace and normalized to have unit norm) $\tilde{h}_{k}$. For each example $x_{i}\rightarrow y_{i}$, the inner product peaks for $k=y_{i}-x_{i}$, indicating that the model extracts the information of $y_{i}-x_{i}$ from the corresponding example $x_{i}\rightarrow y_{i}$.
5.3 Signal Correlation among Examples
@@
Having studied the signal extracted from each individual example in the previous subsection, we next study how signals from different examples interact to execute the ICL task. To do so, we compute the correlation between the extracted signals from different demonstration examples: for each $y_{i}$ token, we first compute the inner product between the residual stream input to head 1, $z_{t}$, and the corresponding task vector $h_{k}$, where $k=y_{i}-x_{i}$. Then, we compute the correlation of these measures across each pair of five positions over $100$ add-k prompts, yielding $\binom{5}{2}$ correlation values per task.
@@
To analyze the correlation, we sum the negative correlation values and positive correlation values respectively for each task, and calculate the various statistics (max, average and min) over all tasks (Table 2). The negative correlation sum is significantly higher than the positive correlation sum for all three statistics, indicating that the signals from any two demonstration examples are mostly negatively correlated. This suggests a self-correction mechanism: intuitively, when the head extracts a noisy signal from one example, signals from subsequent examples are more likely to correct the error, thereby stabilizing the final representation.
Statistics over tasks Sum of $\mathbf{negative}$ correlations Sum of $\mathbf{positive}$ correlations
Average $-2.01$ $0.27$
Min $-1.40$ $0.07$
Max $-2.34$ $0.54$
Table 2: The statistics (min, max, and average) of the absolute values of negative correlation sum and positive correlation sum respectively over the $30$ tasks. The negative correlation sum is significantly higher than the positive correlation sum for all three statistics, which indicates that the signals from any two demonstration examples are mostly negatively correlated, suggesting a self-correction mechanism.
@@ 6 Related Work
Our work builds on a growing body of research that aims to uncover circuits and internal computations of language models (Olah et al., 2020; Wang et al., 2022; Wu et al., 2023; Cunningham et al., 2023; Feng & Steinhardt, 2023; Gould et al., 2023; Park et al., 2024). Below we discuss three particular threads that are most relevant to this paper.
@@
A recent line of work examines how LLMs perform arithmetic (Stolfo et al., 2023; Hanna et al., 2023; Nikankin et al., 2024; Maltoni & Ferrara, 2024), and in particular addition (Nanda et al., 2023; Zhong et al., 2023; Zhou et al., 2024; Kantamneni & Tegmark, 2025). Zhou et al. (2024) find that pre-trained LLMs perform addition using Fourier features, and Kantamneni & Tegmark (2025) find that mid-sized LLMs compute addition using a “clock” algorithm via a helix representation of numbers. Different from them, our work is the first to analyze addition for in-context learning in LLM, and interestingly we find similar representation structures to Zhou et al. (2024) and Kantamneni & Tegmark (2025).
@@
Researchers have constructed detailed models of in-context learning (ICL) for small transformer models in standard supervised learning problems such as linear regression (Garg et al., 2022; Akyürek et al., 2023; Zhang et al., 2023; Li et al., 2023; Wu et al., 2024), as well as more complex settings (von Oswald et al., 2023; Bai et al., 2023; Bietti et al., 2023; Reddy, 2023; Guo et al., 2023; Nichani et al., 2024). For large pretrained models, there exist coarser-grained treatments attributing ICL performance to either induction heads (Olsson et al., 2022; Singh et al., 2024; Crosbie & Shutova, 2025; Bansal et al., 2023) or function vector (FV) heads (Todd et al., 2024; Hendel et al., 2023). Yin & Steinhardt (2025) compares the two types of heads and finds that few-shot ICL performance depends primarily on FV heads. Motivated by this, we study function vector heads in details for a family of ICL tasks, introducing a novel optimization method, which achieves better performance than the method in Todd et al. (2024). Another difference from Todd et al. (2024) is that our tasks have the same input domain, ensuring that the ICL prompts for different tasks differ only in the task information.

### v3

@@ 3.2 Further Refinement via Ablations
We suspect many heads are primarily responsible for storing formatting information (such as ensuring the output appears as a number) rather than encoding information about $k$ itself. Intuitively, while we require the overall signal transmitted by these heads, we do not need any information about the specific value of $k$. To test this hypothesis, we perform mean-ablations: replacing each task-specific signal $h_{k}$ with the overall mean $\overline{h}$ across all values of kk2.2).
@@
$v_{k}=\sum_{h\in\mathcal{H}_{0}}\overline{h}+\sum_{h\in\mathcal{H}_{\text{sig}}\setminus\mathcal{H}_{0}}h_{k}.$ (2)
@@
Focusing on two layers via layer-wise ablation. To efficiently narrow down the important heads, we first perform mean-ablations at the level of layers. From Figure 3(a), we observe that the significant heads $\mathcal{H}_{\text{sig}}$ are concentrated primarily in the middle and late layers. We speculate that heads in the late layers mainly contribute to formatting the output, as they appear too late in the computation to meaningfully interact with the query. After trying different sets of layers, we found that mean-ablating all significant heads outside layers $13$ and $15$ still achieves an intervention accuracy of $0.83$, while mean-ablating any other combination of layers causes negligible drops in accuracy (Appendix B.1, Table 3). After these ablations, only $11$ heads located in layers $13$ and $15$ remain.
Identifying three final heads via head-ablation. To understand the individual contributions of each head within layers 13 and 15, we perform mean-ablations at the level of individual heads. We first assess the intervention accuracy when retaining only the output of a single head while mean-ablating all other significant heads; however, this generally results in low accuracy. We hypothesize that the output magnitude of a single head is too small to significantly influence the model output, even if it encodes task-relevant information. To amplify each head’s effect, we scale its output by a coefficient (e.g., 5). We find that three heads—head 1 = $(15,2)$, head 2 = $(15,1)$, and head 3 = (13,6)(13,6)—achieve intervention accuracies close to the clean accuracy when appropriately scaled, while all other heads show much lower accuracies regardless of scaling (Table 1). This suggests that these three heads individually encode the task information much better than any others.
@@
$(15,2)\text{:=Head 1}$ 66   0.85
$(15,1)\text{:=Head 2}$ 55   0.83
$(13,6)\text{:=Head 3}$ 55   0.66
Any other head  Optimala   0.19
  • a
    Coefficient chosen to maximize accuracy for each head (scanned over integer values).
@@ 4 Characterizing the Aggregator Subspace
For the model to perform add-k, it has to infer the task information (the number $k$) from the ICL demonstrations. Our next goal is thus to understand how task information is represented in the activation space. Since we have identified three aggregator heads that carry almost all of this information, we can now focus on analyzing the representation space of these three heads.
In this section, we dissect their activation spaces in three stages: (1) Localize a six–dimensional task subspace in each head via principal component analysis (PCA) (§4.1); (2) Rotate this subspace into orthogonal feature directions aligned with sinusoidal patterns across kk4.2); (3) Decompose the six-dimensional space into a four-dimensional unit-digit subspace and a two-dimensional magnitude subspace that separately encode the units and tens of the answer (§4.3).
@@ 4.1 Localizing to Six-dimensional Subspace
To reduce the $128$-dimensional head activation to a more tractable space to study, we first perform PCA on the $30$ task vectors and find that just six directions can explain $97\%$ of the task variance (Appendix C, Figure 4(a)). We then check that the function vectors found earlier remain effective after projecting onto the subspace. Specifically, we replace each head vector $h_{k}$ with its projection onto the subspace $\tilde{h}_{k}$ to obtain a new function vector $\tilde{v}_{k}=\sum_{h\in\mathcal{H}}\tilde{h}_{k}$ (a variant of $v_{k}=\sum_{h\in\mathcal{H}}h_{k}$ in §2.2). We find that $\tilde{v}_{k}$ has intervention accuracy $0.76$, which is close to the intervention accuracy of $0.79$ before projection (§3.2). Thus, we confine our study to the concise six-dimensional subspace for each head.
@@ 4.2 Identifying Feature Directions Encoding Periodic Patterns
To understand how the six-dimensional subspace of each head represents task information, we first examine coordinates of head vectors (their inner products with principal components (PC)) as functions of $k$. This reveals partially periodic patterns in the first five PCs (Appendix D.1, Figure 5).
@@
To find trigonometric functions to fit, we searched over different periods and phases and performed least squares regression. We found six trigonometric functions at periods $2$, 55, $10$, $10$, $25$, and $50$ that could be expressed as functions of the top 6 PCs with low regression error (Appendix D.1, Figure 6). We apply the resulting linear transformation to the six PCs to obtain a new set of directions that encode these six pure periodic patterns, which we call feature directions.
@@ 4.3 Decomposing to Subspaces Encoding Subsignals
By construction, the coordinate function of each feature direction (viewed as linear projections of $h_{k}$) is a periodic function of $k$. Mathematically, a feature direction with period $T$ carries task information from the head vectors with “modulo TT”. Based on this, we hypothesize: (i) the feature direction corresponding to period two, which we call the “parity direction”, encodes the parity of $k$ in add-k task; (ii) the subspace spanned by the feature directions with periods $2,5,10$, which we call the “unit subspace”, encodes the unit digit of $k$; (iii) the subspace spanned by the directions with periods $25,50$, which we call the “magnitude subspace”, encodes the coarse magnitude (i.e., the tens digit) of $k$.
We verify these hypotheses through causal intervention. We establish (1) sufficiency by showing that projecting a head vector onto the subspace preserves the relevant task signal; and (2) necessity by showing that projecting a head vector out of the subspace (i.e., onto its orthogonal complement) destroys the relevant task signal. We defer experimental results in Appendix E.
5 Signal Extractors of ICL Demonstrations
Previously, we localized the behavior to three heads and their six-dimensional subspaces, then examined how the model represents the task information within these subspaces. Now, we analyze how the model extracts the task information from the ICL demonstrations.
In this section, we find that: (1) the signal is primarily gathered from the label tokens in each demonstration; (2) each demonstration $x_{i}\rightarrow y_{i}$ individually contributes a signal $y_{i}-x_{i}$ in the subspace even on “mixed” in-context demonstrations with conflicting task information; and (3) when all demonstrations $x_{i}\rightarrow y_{i}$ share the same value for $y_{i}-x_{i}$, the extracted signals exhibit a self-correction behavior.
@@ 5.1 Mathematical Observation: Tracing Subspace back to Previous Tokens
Formally, a head $h$’s output at the last token of a prompt $p$ can be written as $h(p)\;=\;\sum_{t\in p}\alpha_{t}\,O_{h}V_{h}\,z_{t},$
@@
Let $W_{h}$ denote the projection matrix onto the six-dimensional subspace for head $h$. Then the projected signal at the final token, $W_{h}\cdot h(p)$, can be decomposed into contributions from previous tokens as $W_{h}\cdot\alpha_{t}O_{h}V_{h}z_{t}$, each lying in the image of the head subspace under $W_{h}O_{h}V_{h}$. In the following subsection, we analyze the magnitudes and directions of these signals, and study how signals from different demonstrations interact.
5.2 Signal Extractor for each Demonstration
@@
To understand how demonstrations contribute to model generation at the last token, we identify which tokens contribute the most, then examine what information they provide. By the analysis above, the task-signal contribution of each previous token to the final token through the head $h$ is $\alpha_{t}W_{h}O_{h}V_{h}z_{t}$. This can be decomposed into two parts: (1) extracted information: $W_{h}O_{h}V_{h}z_{t}$, the residual stream input projected into the relevant subspace; and (2) aggregation weight: $\alpha_{t}$, the attention score of the final token to the previous token. We plot the norms of the extracted information and the aggregation weights for a random mixed-$k$ ICL prompt (with conflicting $k_{i}$ values), in Appendix F.1. Both the strength of the extracted information and the aggregation weights peak at $y_{i}$ tokens.
@@
We next examine what specific information is extracted from each of these tokens. To do so, we measure the inner product $\langle W_{h}O_{h}V_{h}z_{t},\tilde{h}_{k}\rangle$ between the extracted information and the head vector (projected onto the subspace and normalized to have unit norm) for each task $k$. In Figure LABEL:fig:15_2_extractor_direction, we plot this quantity for a random mixed-$k$ ICL prompt for each token $y_{i}$ ($i\in\{1,\ldots,5\}$) and each task $k\in\{1,\ldots,30\}$. We find that the inner product consistently peaks at $k=y_{i}-x_{i}$, indicating that the model extracts the information of $y_{i}-x_{i}$ from the corresponding demonstration $x_{i}\rightarrow y_{i}$.
5.3 Signal Correlation among Demonstrations
@@
Having studied the signal extracted from each demonstration in the previous subsection, we next study how signals from different demonstrations interact to execute ICL task. To do so, we compute the correlation between the extracted signal from different demonstrations: for each $y_{i}$ token, we first compute the inner product between the residual stream input to head 1, $z_{t}$, and the corresponding task vector $h_{k}$, where $k=y_{i}-x_{i}$. Then, we compute the correlation of these measures across each pair of five positions over $100$ add-k prompts, yielding $\binom{5}{2}$ correlations per task.
@@
To analyze the correlation, we sum the negative correlation values and positive correlation values respectively for each task, and calculate the various statistics (max, average and min) over all tasks (Appendix G.1, Table 8). The negative correlation sum is significantly higher than the positive correlation sum for all three statistics, indicating that the signals from any two demonstrations are mostly negatively correlated. This suggests a self-correction mechanism: intuitively, when the head extracts a noisy signal from one demonstration, signals from subsequent demonstrations are more likely to correct the error, thereby stabilizing the final representation.
@@ 6 Related Work
Our work builds on a growing body of interpretability research that aims to uncover circuits and internal computations of language models. Many studies focus on synthetic tasks or models specifically trained on that task (Nanda et al., 2023; Bietti et al., 2023; Reddy, 2023; Singh et al., 2024). Going beyond to large pretrained language models, some papers study general LLMs and tasks but provide only coarser-grained analyses (Todd et al., 2024; Olsson et al., 2022; Hendel et al., 2023), while others focus on particular model families and specific task classes to obtain more fine-grained insights (Hanna et al., 2023; Feng & Steinhardt, 2023; Wu et al., 2023; Zhou et al., 2024; Panickssery et al., 2024). Our work follows the latter trajectory: we analyze Llama-3 models and the Qwen-2.5 model on a structured set of addition ICL tasks, and we provide a deeper and more detailed account of ICL mechanisms than prior studies of ICL in LLMs (Todd et al., 2024; Olsson et al., 2022; Hendel et al., 2023). Below we discuss three particular threads that are most relevant to this paper.
@@
A recent line of work examines how LLMs perform arithmetic (Stolfo et al., 2023; Hanna et al., 2023; Nikankin et al., 2024; Maltoni & Ferrara, 2024), and in particular addition (Nanda et al., 2023; Zhong et al., 2023; Zhou et al., 2024; Kantamneni & Tegmark, 2025). Zhou et al. (2024) find that pre-trained LLMs perform addition using Fourier features, and Kantamneni & Tegmark (2025) find that mid-sized LLMs compute addition using a “clock” algorithm via a helix representation of numbers. Unlike prior work, we analyze addition in the ICL setting for LLMs, and interestingly we find similar representation structures to them.
@@
Researchers have constructed detailed models of in-context learning (ICL) for small transformer models in standard supervised learning problems such as linear regression (Garg et al., 2022; Akyürek et al., 2023; Zhang et al., 2023; Li et al., 2023; Wu et al., 2024), as well as more complex settings (von Oswald et al., 2023; Bai et al., 2023; Bietti et al., 2023; Reddy, 2023; Guo et al., 2023; Nichani et al., 2024). For large pretrained models, there exist coarser-grained treatments attributing ICL performance to either induction heads (Olsson et al., 2022; Singh et al., 2024; Crosbie & Shutova, 2025; Bansal et al., 2023) or function vector (FV) heads (Todd et al., 2024; Hendel et al., 2023). Yin & Steinhardt (2025) compares the two types of heads and finds that few-shot ICL performance depends primarily on FV heads. Motivated by this, we study function vector heads in detail for a family of few-shot ICL tasks, introducing a novel optimization method, which achieves better performance than the method in Todd et al. (2024). Another difference from Todd et al. (2024) is that our tasks have the same input domain, ensuring that the ICL prompts for different tasks differ only in the task information, which allows for a clearer understanding of ICL mechanism.